OpenStudy (anonymous):

express the confidence interval 0.111

OpenStudy (anonymous):

\(0.333-0.111=0.222,0.222/2=0.111,0.111+0.111=0.222\) hence \(p=0.222\pm0.111\)

OpenStudy (anonymous):

thank you